OpenStudy (anonymous):

Let f be the function f(x)=2xe^x. The graph of f is concave down when a. x<-2 b.x >-2 c. x<-1 d. x>-1 e. x<0

OpenStudy (kinggeorge):

First, you need to take the second derivative. It's a little bit of work, but you get the equation \[2 e^x (2+x)\]Now we want this to be less than 0.

OpenStudy (anonymous):

can you showme howto do the whole problem?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

x<-2

OpenStudy (anonymous):

\[f(x)=2xe^x\] product rule \[f'(x)=2xe^x+2e^x=2(xe^x+e^x)\] second derivative product rule again \[f''(x)=2(xe^x+e^x+e^x)=2(xe^x+2e^x)=2e^x(x+2)\]

OpenStudy (anonymous):

you want to know where the second derivative is negative. \(e^x>0\) so we can ignore that factor and solve \(x+2<0\) making \(x<2\) that is where the second derivative if negative, and so that is where the function is concave down

OpenStudy (anonymous):

oops!

OpenStudy (anonymous):

\[x+2<0\implies x<-2\] must be getting late

OpenStudy (kinggeorge):

Had to step away for a bit. Sorry about that. But now we need to notice that \(2e^x\) is always greater than 0. Thus, if we want the second derivative must be negative, we do what satellite just did above me. So \[x<-2\]